OpenStudy (anonymous):

solve the system x-3=2y and 4x-5y=9. what is the value of x in the solution? A.-2 B.-1 C. 1 D.2

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

No problem any more questions? :)

OpenStudy (uri):

x-3=2y 4x-5y=9 Arrange eq 1 first x-3=2y x-2y=3(multiply by 5) 4x-5y=9(multiply by -2) 5x-10y=15 -8x+10y=-18 ========== -3x=-3 x=-3/-3 x=1 @Daniellee it's your last warning,If i see you giving answers.I will have to talk to mods.

OpenStudy (anonymous):

just a few but i have to go to work, ill be back on arpund 4:00 id that is convenient for you?

OpenStudy (anonymous):

yea just mention me in the questions and @uri ill make sure to explain next time

OpenStudy (anonymous):

ill do that, thank you again(:

OpenStudy (uri):

http://openstudy.com/users/alex_richard_pankey#/updates/51ffcb50e4b0cc46c14b35de You did the same after i told you here.

OpenStudy (anonymous):

no problem! :)

OpenStudy (anonymous):

no you said "now mind explaining? and i just thought hey i already gave the answer soo yea i didnt know you had to explain i will explain next time

OpenStudy (anonymous):

the english section?

OpenStudy (anonymous):

??